OpenStudy (anonymous):

How do I do this? 5X+21=51

OpenStudy (vijay):

6

jimthompson5910 (jim_thompson5910):

5X+21=51 5X=51 -21 5X=30 X = 30/5 X = 6

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

good answer, jim